OpenStudy (anonymous):

(x-8)^2=20 give exact solutions and to three decimal places

OpenStudy (anonymous):

first foil the left hand side

OpenStudy (anonymous):

then subtract 20 from both sides so you get a quadratic equal to zero. then use the quad formula to solve for x.

OpenStudy (anonymous):

(x-4)(x+4)?

OpenStudy (anonymous):

no. (x-8)^2 = (x-8)(x-8) now foil

OpenStudy (anonymous):

x^2-8x-8x-64 x^2-64

OpenStudy (anonymous):

combine the -8x and -8x to get -16x, and (-8)*(-8) = +64, so after foiling you should get x^2 -16x +64

OpenStudy (anonymous):

oh dang so I add instead of subtract them?

OpenStudy (anonymous):

yes, because they are both negative.

OpenStudy (anonymous):

ok so how do i get my solutions?